When an electric current is passed through water, hydrogen and oxygen gases are liberated. What STP volume of oxygen gas is obtained from the electrolysis of 60.0 mL of water?
Express your answer with the appropriate units.
1
Expert's answer
2020-12-01T09:31:20-0500

Volume of water = 60.0ml

Density of water at 25°C= 1.000g/ml

Mass of water = density x volume

Mass of water = 1.000 x 60 = 60g

Molar mass of water = 18g/mol

Number of moles of water = mass/molar mass = 60/18= 3.33mol

From the half equation

2H2O ------> O2 + 4H+ + 4e- (oxidation)

2 moles of water produces 1 mole of oxygen gas at STP

3.33 moles of water will therefore produce 1/2x 3.33 = 1.67 moles of O2 at STP


At STP, one mole of O2 occupies 22.4L

1.67mol will therefore occupy 1.67 x 22.4 = 37.41Litres

Therefore, the volume of oxygen gas obtained from the electrolysis is 37.41L.